# On the Convergence of Muon and Beyond

### v2

@@ 6.2 Analysis of Muon
Recent studies have elucidated the underlying mechanisms and convergence properties of the Muon optimizer from multiple theoretical perspectives. A central viewpoint reinterprets modern optimizers as steepest descent methods under norm constraints. Although this analysis does not directly incorporate momentum, it offers valuable insights into the update directions of optimizers like Muon and Adam Bernstein & Newhouse (2024). Building on this, subsequent research has established a theoretical link between the Muon algorithm with weight decay and the stochastic Frank-Wolfe method, specifically identifying Muon as an instance of stochastic Frank-Wolfe under a spectral norm constraint Sfyraki & Wang (2025). This provides a foundational understanding of its behavior from a constrained optimization perspective. Other work has characterized Muon as a special case of Lion-$\mathcal{K}$. Separately, its convergence to a Karush-Kuhn-Tucker (KKT) point has been proven in both deterministic and stochastic gradient settings Chen et al. (2025). These theories, together with research on stochastic conditional gradient methods from a norm-constrained viewpoint, form the theoretical underpinnings for Muon Pethick et al. (2025). Beyond its theoretical framework, Muon’s implicit bias has garnered significant attention. It has been shown that Muon tends to converge to a max-margin solution with respect to the spectral norm of the classifier’s weight matrix, highlighting a distinct difference in implicit regularization and generalization preference compared to Adam Fan et al. (2025). A unifying preconditioning perspective explains this phenomenon through matrix polar decomposition, which explicitly distinguishes between curvature anisotropy and gradient anisotropy in optimization problems, thereby clarifying the behavioral differences between Muon and Adam when handling different types of parameters Lau et al. (2025). Regarding convergence analysis, while rigorous proofs and convergence rates for Muon in non-convex settings have been provided under various smoothness assumptions (e.g., Frobenius-norm or spectral-norm Lipschitz smoothness) Shen et al. (2025), the validity of some early analyses has been questioned. For instance, certain convergence results are considered dubious due to the use of incorrect mathematical inequalities in their proofs Li & Hong (2025). Furthermore, other analyses, despite proving the convergence of several Muon variants, have indicated slow convergence rates or the need for stringent conditions to reach a desired stationary point Sato et al. (2025). On a practical level, particularly in the pre-training of large-scale language models (LLMs), weight decay has been shown to be an indispensable component. Concurrently, methods for effectively transferring learning rates from Adam to Muon, based on RMS analysis, have been proposed to enhance its scalability and utility Liu et al. (2025a).
@@ 7 Conclusion
In this work, we established a rigorous theoretical foundation for the Muon optimizer, addressing the significant gap between its empirical success and its formal analysis. We introduced Muon-MVR2, a variance-reduced variant, and provided the first proof that it achieves an optimal convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$ in stochastic non-convex settings, matching the theoretical lower bound Arjevani et al. (2023). Furthermore, we demonstrated that under the Polyak-Łojasiewicz condition, the Muon-MVR1, and Muon-MVR2 variants all attain sublinear non-ergodic convergence. Our theoretical results are validated by extensive experiments on CIFAR-10 and C4 benchmarks, which confirm the practical acceleration and superior performance of Muon-MVR2 over its standard counterpart and other widely-used optimizers. This research solidifies the Muon framework by providing robust theoretical guarantees and a practically effective new variant for training deep learning models.
@@
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta^{2}n}{2}$
@@
$\leq f(\mathbf{X}_{t})-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle-\eta_{t}\langle\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
@@
$\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F}$
@@
$\leq\eta_{t}\left(\frac{\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{n}{2\alpha}\right)$
@@
$\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}\leq\frac{L\eta_{t}^{2}n}{2}$
@@
$\quad+2\beta_{t+1}(1-\beta_{t+1})\langle\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1}),\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}$
@@
$+2\beta_{t+1}^{2}\mathbb{E}[\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle]$
@@
$\quad+(1-\beta_{t+1})^{2}\mathbb{E}\left[\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$
@@
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}t\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}$ (6)
@@
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}E_{t}\cdot(t+1)^{-p}\cdot(2-c),$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}0$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}^{2/3}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+Ln\eta_{t}^{4/3}$
@@
$+\underbrace{\frac{L\eta_{t}^{4/3}}{2}+\left(\frac{\eta_{t}^{2/3}}{2L}+\frac{\eta_{t}^{2/3}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{4/3}}_{\Gamma_{t}}$
@@
$\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{L^{2}\eta_{t}^{2}n}{1-\beta_{t+1}}+(1-\beta_{t+1})^{2}\sigma^{2}$
@@
$\frac{\eta_{t}^{2}}{1-\beta_{t+1}}=\frac{t^{-3/2}}{(t+1)^{-1/2}}=\frac{\sqrt{t+1}}{t^{3/2}}\leq\frac{2\sqrt{2}}{t+1}$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\alpha_{t+1}^{2}(2\sqrt{2}L^{2}n+\sigma^{2})$
@@
$\alpha_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1})$
@@
$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{1}=0)$
@@
$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)$
@@
$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}$
@@
$\mathbf{M}_{t}$ $=\beta_{t}\mathbf{C}_{t}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})$
@@
$=(1-\alpha_{t})\mathbb{E}\|\mathbf{C}_{t-1}-\nabla f(\mathbf{X}_{t-1})\|_{F}^{2}+\alpha_{t}^{2}(2\sqrt{2}Ln+\sigma^{2})$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+A_{t+1}$
@@
$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{0}=0)$
@@
$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)$
@@
$=\|\beta_{t}(\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t}))+(1-\beta_{t})(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t}))\|_{F}$
@@
$\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\leq 2\beta_{t}^{2}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t})\|_{F}^{2}$
@@
$\leq 2\mathbb{E}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\sigma^{2}$
@@
$=6\sigma^{2}+4(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)$
@@
$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}^{\prime}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}$
@@
$\stackrel{{\scriptstyle(\bullet)}}{{\geq}}\ln(t)t^{-q}\cdot\left\{\sqrt{\frac{1}{2}}-\frac{1}{4}\right\}\geq 0$
Here, $(\circ)$ holds because $\ln(t+1)>\ln(t)$ and $(t+1)^{q}>t^{q}$ for all $t\geq 1$ and $q\in(0,1/2]$. The inequality $(\star)$ holds because $\frac{\ln(t+1)}{\ln(t)}>1$ for $t>1$, and the function $(\frac{t}{t+1})^{q}$ is decreasing in $q$, thus its minimum on $(0,1/2]$ is achieved at $q=1/2$. The final inequality $(\bullet)$ holds because $t\geq 1$ implies $\sqrt{\frac{t}{t+1}}\geq\sqrt{\frac{1}{2}}$, and $\sqrt{\frac{1}{2}}>1/4$.
@@
$\stackrel{{\scriptstyle(\bullet)}}{{\leq}}0$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}+P_{t+1}$
@@
$P_{t+1}$ $=\mathbb{E}\|\Delta_{t}\|_{F}^{2}(\beta_{t+1}(1-\gamma_{t+1})-A_{t+1})^{2}-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}$
@@
$=(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})+\beta_{t+1}(\mathbf{M}_{t}+\gamma_{t+1}\Delta_{t})$
@@
$=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}(\mathbf{S}_{t}+\delta_{t}+\gamma_{t+1}\Delta_{t})$
@@
$\delta_{t}=\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})=-\mathbb{E}_{t+1}[\Delta_{t}]$
@@
$\mathbf{S}_{t+1}=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\gamma_{t+1}\Delta_{t}-\mathbb{E}\Delta_{t})$
@@
$\quad+\underbrace{2\beta_{t+1}(\gamma_{t+1}-1)\mathbb{E}\langle\Delta_{t},(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\rangle}_{\text{Term A.3}}$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\mathbb{E}\|\mathbf{R}_{t+1}\|_{F}^{2}+2\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}-\mathbb{E}\Delta_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\sigma^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+P_{t+1}$
@@
$\gamma_{t+1}=1-\frac{A_{t+1}}{\beta_{t+1}}\text{ or }\gamma_{t+1}=\gamma=1$ (16)
@@
$P_{t+1}=-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}\leq 0\text{ or }P_{t+1}=0$
@@
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}$

### v3

@@ 5.2 Analysis of Muon
Recent theoretical analyses have clarified the mechanisms and convergence of the Muon optimizer. One central perspective frames modern optimizers as steepest descent methods under specific norm constraints; while this view does not directly account for momentum, it offers crucial insights into the update directions of optimizers like Muon and Adam bernstein2024old. Building on this, subsequent work has linked Muon with weight decay to the stochastic Frank-Wolfe method, identifying it as an instance operating under a spectral norm constraint sfyraki2025lions. From a constrained optimization viewpoint, Muon has also been characterized as a special case of Lion-$\mathcal{K}$, with its convergence to a KKT point proven in both deterministic and stochastic settings chen2025muon. These analyses, alongside broader research on norm-constrained stochastic conditional gradient methods, form Muon’s theoretical underpinnings pethick2025training. Beyond its formal framework, Muon’s implicit bias has become a key area of investigation. It has been demonstrated that Muon tends to converge toward max-margin solutions with respect to the spectral norm of the weight matrix, revealing a distinct implicit regularization and generalization preference compared to Adam fan2025implicit. A unifying preconditioning perspective, based on matrix polar decomposition, explains this behavior by distinguishing between curvature and gradient anisotropy. This decomposition clarifies the differing ways Muon and Adam handle various parameter types lau2025polargrad. The convergence analysis of Muon has also seen active development. Although rigorous proofs and iteration complexity in non-convex settings have been established under various smoothness assumptions shen2025convergence, the validity of some early results has been challenged. For instance, certain proofs were found to rely on incorrect mathematical inequalities, casting doubt on their conclusions Li2025ANO. Other analyses, while confirming the convergence of several Muon variants, report slow convergence rates or the need for stringent conditions to reach a stationary point Sato2025ConvergenceBA. On a practical level, especially for pre-training large-scale language models (LLMs), weight decay has been identified as an indispensable component for Muon. To enhance its scalability and utility, methods based on RMS analysis have also been proposed to effectively transfer learning rates from Adam to Muon Liu2025MuonIS.
@@ 6 Conclusion
In this work, we establish a rigorous theoretical foundation for the Muon optimizer, addressing the gap between its empirical success and formal analysis. We analyze two momentum-based variance-reduced variants of Muon: a one-batch version (Muon-MVR1) and a two-batch version (Muon-MVR2). We provide the first proof that Muon-MVR2 achieves the optimal iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$ in stochastic nonconvex settings, matching the theoretical lower bound arjevani2023lower. Furthermore, we show that under the Polyak–Łojasiewicz condition, Muon-MVR1 and Muon-MVR2 attain sublinear nonergodic convergence rates of $\tilde{\mathcal{O}}(T^{-1/2})$ and $\tilde{\mathcal{O}}(T^{-2/3})$, respectively. Our theoretical results are validated by extensive experiments on the CIFAR-10 and C4 benchmarks, which confirm the practical acceleration and superior performance of Muon-MVR2 over its standard counterpart and other widely used optimizers. Overall, this research strengthens the Muon framework by providing robust theoretical guarantees and a practically effective new variant for deep learning training.
@@
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta^{2}n}{2}.$
@@
$\leq f(\mathbf{X}_{t})-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle-\eta_{t}\langle\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}.$
@@
$\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F}.$
@@
$\leq\eta_{t}\left(\frac{\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{n}{2\alpha}\right).$
@@
$\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}\leq\frac{L\eta_{t}^{2}n}{2}.$
@@
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta_{t}^{2}n}{2}.$
@@
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$
@@
$\quad+2\beta_{t+1}(1-\beta_{t+1})\langle\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1}),\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}.$
@@
$+2\beta_{t+1}^{2}\mathbb{E}[\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle].$
@@
$\quad+(1-\beta_{t+1})^{2}\mathbb{E}\left[\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right].$
@@
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$ (6)
@@
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}E_{t}\cdot(t+1)^{-p}\cdot(2-c)$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}0,$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}^{2/3}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+Ln\eta_{t}^{4/3},$
@@
$+\underbrace{\frac{L\eta_{t}^{4/3}}{2}+\left(\frac{\eta_{t}^{2/3}}{2L}+\frac{\eta_{t}^{2/3}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{4/3}}_{\Gamma_{t}},$
@@
$\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{L^{2}\eta_{t}^{2}n}{1-\beta_{t+1}}+(1-\beta_{t+1})^{2}\sigma^{2}.$
@@
$\frac{\eta_{t}^{2}}{1-\beta_{t+1}}=\frac{t^{-3/2}}{(t+1)^{-1/2}}=\frac{\sqrt{t+1}}{t^{3/2}}\leq\frac{2\sqrt{2}}{t+1}.$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\alpha_{t+1}^{2}(2\sqrt{2}L^{2}n+\sigma^{2}).$
@@
$\alpha_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1}).$
@@
$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{1}=0).$
@@
$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1).$
@@
$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}.$
@@
$\mathbf{M}_{t}$ $=\beta_{t}\mathbf{C}_{t}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t}).$
@@
$=(1-\alpha_{t})\mathbb{E}\|\mathbf{C}_{t-1}-\nabla f(\mathbf{X}_{t-1})\|_{F}^{2}+\alpha_{t}^{2}(2\sqrt{2}Ln+\sigma^{2}),$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+A_{t+1}.$
@@
$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{0}=0).$
@@
$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1).$
@@
$=\|\beta_{t}(\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t}))+(1-\beta_{t})(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t}))\|_{F}.$
@@
$\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\leq 2\beta_{t}^{2}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t})\|_{F}^{2}.$
@@
$\leq 2\mathbb{E}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\sigma^{2}.$
@@
$=6\sigma^{2}+4(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1),$
@@
$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}^{\prime}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}.$
@@
$\stackrel{{\scriptstyle(\bullet)}}{{\geq}}\ln(t)t^{-q}\cdot\left\{\sqrt{\frac{1}{2}}-\frac{1}{4}\right\}\geq 0,$
where $(\circ)$ holds because $\ln(t+1)>\ln(t)$ and $(t+1)^{q}>t^{q}$ for all $t\geq 1$ and $q\in(0,1/2]$. The inequality $(\star)$ holds because $\frac{\ln(t+1)}{\ln(t)}>1$ for $t>1$, and the function $(\frac{t}{t+1})^{q}$ is decreasing in $q$, thus its minimum on $(0,1/2]$ is achieved at $q=1/2$. The final inequality $(\bullet)$ holds because $t\geq 1$ implies $\sqrt{\frac{t}{t+1}}\geq\sqrt{\frac{1}{2}}$, and $\sqrt{\frac{1}{2}}>1/4$.
@@
$\stackrel{{\scriptstyle(\bullet)}}{{\leq}}0.$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}+P_{t+1},$
@@
$P_{t+1}$ $=\mathbb{E}\|\Delta_{t}\|_{F}^{2}(\beta_{t+1}(1-\gamma_{t+1})-A_{t+1})^{2}-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}.$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}.$
@@
$=(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})+\beta_{t+1}(\mathbf{M}_{t}+\gamma_{t+1}\Delta_{t}).$
@@
$=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}(\mathbf{S}_{t}+\delta_{t}+\gamma_{t+1}\Delta_{t}).$
@@
$\delta_{t}=\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})=-\mathbb{E}_{t+1}[\Delta_{t}].$
@@
$\mathbf{S}_{t+1}=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\gamma_{t+1}\Delta_{t}-\mathbb{E}\Delta_{t}).$
@@
$\quad+\underbrace{2\beta_{t+1}(\gamma_{t+1}-1)\mathbb{E}\langle\Delta_{t},(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\rangle}_{\text{Term A.3}}.$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\mathbb{E}\|\mathbf{R}_{t+1}\|_{F}^{2}+2\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}-\mathbb{E}\Delta_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},$
@@
$B_{t+1}$ $=(1-\beta_{t+1})\mathbb{E}\langle\Delta_{t},\mathbf{R}_{t+1}\rangle+\beta_{t+1}\mathbb{E}\langle\Delta_{t},\mathbf{S}_{t}\rangle.$
@@
$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\sigma^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+P_{t+1},$
@@
$\gamma_{t+1}=1-\frac{A_{t+1}}{\beta_{t+1}}\text{ or }\gamma_{t+1}=\gamma=1.$ (16)
@@
$P_{t+1}=-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}\leq 0\text{ or }P_{t+1}=0.$
@@
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2},$